Solving this following ODE by Frobenius Method
X^2y’’+(4x-x^2)y’+(2-x)y=0 full solution step-by-step
1
Expert's answer
2019-05-22T09:06:39-0400


"y=\sum_{k=0}^\infty A_k x^{k+r}"

"y'=\sum_{k=0}^\infty A_k(k+r) x^{k+r-1}"

"y''=\sum_{k=0}^\infty A_k (k+r)(k+r-1)x^{k+r-2}"

"x^2\sum_{k=0}^\infty A_k (k+r)(k+r-1)x^{k+r-2}+"

"(4x-x^2)\sum_{k=0}^\infty A_k(k+r) x^{k+r-1}+"

"(2-x)\sum_{k=0}^\infty A_k x^{k+r}="

"\sum_{k=0}^\infty (A_k (k+r)(k+r-1)+4A_k(k+r)+2A_k)x^{k+r}-"

"-\sum_{k=0}^\infty A_k(k+r+1)x^{k+r+1}="

"\sum_{k=0}^\infty A_k((k+r)(k+r-1)+4(k+r)+2)x^{k+r}-"


"-\sum_{k=1}^\infty A_{k-1}(k+r)x^{k+r}="

"[(k+r)(k+r-1)+4(k+r)+2="

"k^2+2kr+r^2-k-r+4(k+r)+2="

"(k+r)^2+3(k+r)+2=(k+r+1)(k+r+2) ]"

"A_0(r+1)(r+2)x^r+\sum_{k=1}^\infty(A_k(k+r+1)(k+r+2)-A_{k-1}(k+r))x^{k+r}"

indicial polynom:

"(r+1)(r+2)=0,r_1=-1,r_2=-2"

"A_k(k+r+1)(k+r+2)-A_{k-1}(k+r)=0"

"A_k(k+r+1)(k+r+2)=A_{k-1}(k+r)""case 1: k=1,r=-2"

"A_1(1-2+1)(1-2+2)=A_0(1-2)"

"A_0=0,A_k=0;""y = 0"


"case 2: k=1,r=-1"

"A_1(1-1+1)(1-1+2)=A_0(1-1) = 0"

"A_0\not = 0 , A_1=0,A_2=0,..."

"y=\frac{A_0}{x}; y = \frac{0}{x}=0, when A_0=0."

Answer: y=A0/x, A0 is real number.